# kNN-CLIP: Retrieval Enables Training-Free Segmentation on Continually Expanding Large Vocabularies

### v1

@@ Abstract
Rapid advancements in continual segmentation have yet to bridge the gap of scaling to large continually expanding vocabularies under compute-constrained scenarios. We discover that traditional continual training leads to catastrophic forgetting under compute constraints, unable to outperform zero-shot segmentation methods. We introduce a novel strategy for semantic and panoptic segmentation with zero forgetting, capable of adapting to continually growing vocabularies without the need for retraining or large memory costs. Our training-free approach, kNN-CLIP, leverages a database of instance embeddings to enable open-vocabulary segmentation approaches to continually expand their vocabulary on any given domain with a single-pass through data, while only storing embeddings minimizing both compute and memory costs. This method achieves state-of-the-art mIoU performance across large-vocabulary semantic and panoptic segmentation datasets. We hope kNN-CLIP represents a step forward in enabling more efficient and adaptable continual segmentation, paving the way for advances in real-world large-vocabulary continual segmentation methods.
Figure 1: We propose kNN-CLIP to continually expand the vocabulary space of segmentation models. Our approach adapts to concept customization and identifying long-tailed concepts like famous personas. The above image is a qualitative demonstration of our work.
@@ 1 Introduction
The holy grail in image segmentation is to accurately segment any concept images based on a text query. This task was made more achievable through the rapid advances of vision-language models (VLMs) like CLIP, which when applied to semantic and panoptic segmentation, showing promise in handling a broad vocabulary of visual data. However, while they excel in common categories on segmentation benchmarks, they fall short in achieving broad-vocabulary segmentation (Shao et al., 2023; Udandarao et al., 2024).
@@
Our analysis begins by identifying that the success of these models on standard datasets stems from their fine-tuning on specific datasets with detailed labels, such as COCO Stuff (Caesar et al., 2018), which often overlaps with the classes of many other benchmarks with fine-grained annotations. For example, COCO-Stuff and ADE-20K have a class overlap of 73 out of 150 classes. Previous works (Sun et al., 2023) have shown that removing this overlapping vocabulary results in poor segmentation performance. We further discover that this finetuning significantly reduces the open-vocabulary segmentation abilities derived from the pre-trained CLIP model, attributable to catastrophic forgetting. This raises the question: Can we enhance these models to a continually-growing vocabulary with data without catastrophic forgetting?
First, we investigate continual training of these VLM segmentation algorithms to data with new categories. We consistently find that continual training of these models causes concept forgetting by encoding new information into a limited, poorly understood parameter space inadvertently overwriting past knowledge. However, recent state-of-the-art open-vocabulary segmentation approaches (Yu et al., 2023; Liu et al., 2023b) still rely on fine-tuning with datasets containing fine-grained annotations like masks and boxes. This limits the effective vocabulary of these open-vocabulary model, i.e., we observe that training significantly narrows down the vocabulary size of these models, restricting their ability to generalize to a wider range of objects and concepts [^1] Concept forgetting is a term borrowed from Mukhoti et al. (2023), who demonstrated a similar degradation of open-vocabulary performance after fine-tuning CLIP models but on classification tasks. The primary question this work tackles is how to expand the knowledge of the model given in-domain data without any training.
Our novel method, kNN-CLIP, circumvents the need for retraining by using a retrieval database that matches images with text descriptions, updating a support set with new data in a single pass without storing any previous images for replay. In contrast to traditional continual learning techniques, we can guarantee that the model never forgets previously seen data, learns with a single pass, optimizes memory by only storing features, and crucially, expands its vocabulary with minimal computational resources due to no additional training.
Our straightforward and effortless approach significantly improves the performance of leading semantic and panoptic segmentation algorithms without the need for additional training. We apply our method to the state-of-the-art open-vocabulary semantic and panoptic segmentation framework FC-CLIP (Yu et al., 2023), and achieve notable increases in performance (mIoU) across various challenging datasets—A-847, PC-459, and A-150—by +2.6, +1.7 and +7.2 points, respectively. This improvement is achieved with small computational and memory overheads, and without any training cost. As shown in Figure 1, we also qualitatively demonstrate that our method can be used for multiple purposes including concept customization and person identification.
Overall, the key contributions of our study include:
  • Identifying Degradation in Open-Vocabulary Segmenters: We uncover the paradox where fine-tuning VLM models for segmentation using downstream annotations dramatically reduces their ability to recognize the broad VLM vocabulary, illustrating that catastrophic forgetting in open-vocabulary restricts the scope of open-vocabulary segmentation.
@@
  • Training-Free Continual Vocabulary Expansion: We introduce a novel technique, kNN-CLIP, that continually expands the vocabulary of image segmentation models without additional training. This method utilizes a continually expanding support set, recasting the problem of concept prediction as an image retrieval task.
@@ 2 Related Works
Retrieval-Augmented Models. In the NLP domain, Retrieval-Augmented Generation (RAG) has been shown as a promising technique for enhancing Large Language Models (LLM) by leveraging external structured data (Guu et al., 2020; Wu et al., 2022; Borgeaud et al., 2022; Liu et al., 2020; Peters et al., 2019; Yu et al., 2021; Izacard et al., 2022; Ram et al., 2023). The dynamic nature of RAG facilitates continuous knowledge updates, enabling models to integrate domain-specific information seamlessly. This is particularly beneficial for knowledge-intensive tasks (Lewis et al., 2020; Wang et al., 2022; Khandelwal et al., 2019; Shi et al., 2023; Petroni et al., 2023).Inspired by its success in NLP, researchers are now exploring the application of RAG in computer vision tasks (Chen et al., 2022a; Marino et al., 2021; Roth et al., 2022; Shen et al., 2022; Sheynin et al., 2022; Yasunaga et al., 2022; Blattmann et al., 2022; Chen et al., 2022b; Liu et al., 2023a; Udandarao et al., 2023; Iscen et al., 2023; Prabhu et al., 2023c; Balazevic et al., 2024). For instance, REACT (Liu et al., 2023a) presents a methodology aimed at retrieving relevant knowledge and learning customized visual modules accordingly for specific domains. Meanwhile, SuS-X (Udandarao et al., 2023) and C2C (Prabhu et al., 2023c) introduce frameworks that utilizes retrieved support sets for far more accurate CLIP classification to open-set domains. Although these two works show promises in applying retrieval-based methods to visual perception, they did not discuss how to apply such techniques successfully in visual segmentation. Hummingbird (Balazevic et al., 2024) suggests employing a straightforward non-parametric nearest neighbor retrieval as the decoder for dense scene understanding tasks. Moreover, beyond integrating new visual features, RECO (Iscen et al., 2023) demonstrates the efficacy of incorporating text representations by cross-modal fusion alongside original and retrieved embeddings. However, both Hummingbird (Balazevic et al., 2024) and RECO (Iscen et al., 2023) require training to better integrate retrieval augmentation into their pipelines; in contrast, our method demands no training efforts while providing faster adaptability and strong performance.
@@
Open Vocabulary Learning in Image Understanding. Benefiting from advancement in vision-language model (Radford et al., 2021; Jia et al., 2021), visual models demonstrate the potential of open vocabulary image understanding to break the constraints of pre-defined closed-set concept. In particular for dense prediction tasks (Yu et al., 2023; Rao et al., 2022; Ding et al., 2022; Chen et al., 2023; Karazija et al., 2023; Liang et al., 2023; Li et al., 2022a; Xu et al., 2023a; b; 2022b; 2022a; Sun et al., 2023; Wang et al., 2023b; Lüddecke & Ecker, 2022; Xie et al., 2023), one line of works such as SimBaseline (Xu et al., 2022b) and OVSeg (Liang et al., 2023) focus on leveraging cross-modality supervision from CLIP to align class-agnostic mask proposal with language concepts. ODISE (Xu et al., 2023a) and FC-CLIP (Yu et al., 2023), further explore such idea with improved visual encoder and mask proposal strategy to achieve promising performance on both semantic and panoptic segmentation. However, a critical limitation of these works is catastrophic forgetting during fine-tuning on concept-restricted dense prediction datasets, leading to severe degradation in model vocabulary size. To address this challenge, training-free method such as ReCO (Shin et al., 2022) and CaR (Sun et al., 2023) aim to achieve efficient adaptation of CLIP model through recurrently pruning the label space. While these method circumvent the problem of catastrophic forgetting, they are still unable to accommodate continually expanding large vocabulary. Thus, in this work, we leverage retrieval augmentation for continual open vocabulary image understanding.
@@
Continual Learning for Semantic Segmentation. As opposed to open vocabulary semantic segmentation, continual learning does not intend to encompass an extremely large vocabulary space at the first place but preserve the ability to expand the vocabulary space continuously. Nevertheless, continual learning suffers from catastrophic forgetting and semantic drift. To alleviate these issues, iCaRL (Rebuffi et al., 2017) proposes to replay the most representative samples during the continual learning stage, and following works attempt to optimize the associated memory burden  (Cha et al., 2021; Zhu et al., 2023; Wang et al., 2023a; Kalb et al., 2022; Kim et al., 2022). Moreover, feature-replay scheme is proposed by ALIFE (Oh et al., 2022) to reduce memory requirements. More works explore how to leverage memorized features to preserve known knowledge  (Yoon et al., 2022; Yu et al., 2020; Wang et al., 2021; Michieli & Zanuttigh, 2021; Lin et al., 2022a). One notable shortcoming of using memorized features is that the representation capacity of features might be insufficient, hindering the performance of continual learning (Yuan & Zhao, 2023). Either using hyper-class knowledge (Shi et al., 2022) or dynamically updating stored features (Liu et al., 2022; Lin et al., 2022b; Li et al., 2022b) mitigates the problem but still not circumventing it inherently. Serving the same purpose of expanding vocabulary space dynamically, our method resembles the feature-replay methods, as we also construct a support set of features. Nonetheless, our method operates without the need for training and utilizes a support set, storing powerful features learned through self-supervised learning techniques (Caron et al., 2021), with low memory requirements.
3 Continually Expanding Retrieval-Augmented Dense Prediction
@@
We introduce a novel, training-free framework for continual dense prediction with expanding vocabularies, applicable on diverse domains and across dense prediction tasks. Inspired by the Retrieval-Augmented Generation (RAG) method used in Large Language Models, our framework leverages a customizable embedding database to incorporate domain-specific knowledge directly at inference, with no additional training.
@@
This approach is applicable to a range of dense prediction tasks including semantic and panoptic image segmentation. We first evaluate the confidence level of CLIP classification results for each query mask. For queries falling below a given confidence threshold, we extract DINO features and use mask pooling to generate query embeddings. These embeddings are then matched against a vectorized database using a kNN search algorithm based on cosine similarities. For the most similar embeddings found, we create a set of confidence-aware pseudo-logits by stacking similarities across class vectors, guided by the labels of these embeddings. The results from this retrieval process are then combined with the initial CLIP results through a given weighting parameter $\lambda$. We now provide details of our training-free method. We first describe the embedding construction process and detail the query mechanisms.
@@ 3.1 Designing Continually Expanding Embedding Databases
One appealing characteristic of our framework is the dynamic extendability of the database and the customization of the database construction based on personalized interest. Depending on the purpose specified by the user-defined applications (fine-grained setting such as smart retail, species classification, etc.), the database could be sourced from various domains through manual curation, online web-based collection, or synthetic data sources.
@@
To obtain distinctive vectorized representations for the database images, we employ mask-pooling on features extracted by pre-trained encoders from the images as shown in Figure 2. Our input consists of an incoming image ($x_{i}\in\mathcal{R}^{3\times h\times w}$) with it’s corresponding $k$ class-agnostic masks, which could be semantic, instance dense masks, or bounding box masks ($M_{i}\in\mathcal{R}^{k\times h\times w}$) and class annotations ($C_{i}$) containing the class $\{c{{}_{ji}}\}_{j=1}^{k}$ corresponding to mask $\{m_{ji}\}_{j=1}^{k}=M_{i}$. We extract features of this image using a Vision Transformer (ViT, Dosovitskiy et al. (2021))-based encoder ($\mathcal{E}$), yielding $h_{i}=\mathcal{E}(x_{i})$. Here, $h_{i}\in\mathcal{R}^{d\times h^{\prime}\times w^{\prime}}$ with $d$ represents the embedding dimension of the feature, and $h^{\prime}$, $w^{\prime}$ represent the dimensions of the feature map. The embedding dimensions $h^{\prime}$, $w^{\prime}$ are determined by image size divided by the patch size of the vision transformer, thus usually shrunken by a magnitude of $10$ each, and efficiently representing images with small feature maps guarantees the memory efficiency of our method. The ground-truth masks $M_{i}$ are resized to $M^{\prime}_{i}\in\mathcal{R}^{k\times h^{\prime}\times w^{\prime}}$ to align with the feature map dimensions. By applying mask average pooling along $h^{\prime}$ and $w^{\prime}$, the final embedding sets $\{e_{ji}\}_{j=1}^{k}$ on image $x_{i}$ is obtained by:
$E_{i}\in\mathcal{R}^{k\times d}=\{e_{ji}=\frac{\sum_{h^{\prime}}\sum_{w^{\prime}}(h_{i}\times m^{\prime}_{ji})}{\sum_{h^{\prime}}\sum_{w^{\prime}}m^{\prime}_{ij}}\}_{j=1}^{k}$ (1)
Then, the vectorized embedding database is constructed by concatenating the embedding sets and class sets from all images, so that $D_{vec}=\{{E_{i},C_{i}}\}_{i=1}^{n}=\{{e_{i^{\prime}},c_{i^{\prime}}}\}_{i^{\prime}=1}^{n^{\prime}}$, if there are $n^{\prime}$ embeddings in total. New embeddings can also be continually added to the database repeating the above process. Note that we do not store any images for replay, enabling low storage-costs and allowing deletion of previously seen data.

### v2

@@ Abstract
Continual segmentation has not yet tackled the challenge of improving open-vocabulary segmentation models with training data for accurate segmentation across large, continually expanding vocabularies. We discover that traditional continual training results in severe catastrophic forgetting, failing to outperform a zero-shot segmentation baseline. We introduce a novel training-free strategy, kNN-CLIP, which augments the model with a database of instance embeddings for semantic and panoptic segmentation that achieves zero forgetting. We demonstrate that kNN-CLIP can adapt to continually growing vocabularies without the need for retraining or large memory costs. kNN-CLIP enables open-vocabulary segmentation methods to expand their vocabularies on any domain with a single pass through the data, while only storing compact embeddings. This approach minimizes both compute and memory costs. kNN-CLIP achieves state-of-the-art performance across large-vocabulary semantic and panoptic segmentation datasets. We hope kNN-CLIP represents a significant step forward in enabling more efficient and adaptable continual segmentation, paving the way for advances in real-world large-vocabulary continual segmentation methods.
Figure 1: We propose kNN-CLIP to continually expand the vocabulary space of segmentation models. Our approach adapts to concept customization and identifying long-tailed concepts, a known challenge for CLIP models (Udandarao et al., 2024). For concept customization, we build the supporting database for each long-tailed concept efficiently using C2C (Prabhu et al., 2023c). We use EntitySeg (Qi et al., 2022) to generate class-agnostic masks for entities in the database and kNN-CLIP to label these masks. At inference time, we filter the masks from EntitySeg based on confidence thresholds for both the mask and class predictions.
@@ 1 Introduction
The ultimate goal in image segmentation is to accurately segment any image according to a text query across open-ended classes. This task was made more achievable through the rapid advances of vision-language models (VLMs) like CLIP (Radford et al., 2021b), which when applied to semantic and panoptic segmentation, shows promise in handling a broad vocabulary of visual data (Zhou et al., 2021; Rao et al., 2022). However, recent state-of-the-art open-vocabulary segmentation approaches (Yu et al., 2023; Liu et al., 2023c) still rely on fine-tuning with datasets containing fine-grained annotations like masks and boxes. Although they excel in common categories on segmentation benchmarks, they fall short of achieving broad-vocabulary segmentation (Li et al., 2023a; Sun et al., 2023b).
@@
We identify that the success of these models on standard datasets is due to fine-tuning on specific datasets with detailed labels, such as COCO-Stuff (Caesar et al., 2018), which often overlap with the classes of many other benchmarks with fine-grained annotations. For example, COCO-Stuff and ADE-20K share 73 out of 150 classes. Crucially, removing this overlapping vocabulary results in poor segmentation performance (Sun et al., 2023a). Hence, the effective vocabulary of these open-vocabulary models is limited to COCO-Stuff, i.e. fine-tuning significantly narrows down the vocabulary size of these models, restricting their ability to generalize to an open-vocabulary of objects and concepts [^1] This phenomenon of degradation of open-vocabulary performance after fine-tuning CLIP models, known as concept forgetting (Mukhoti et al., 2023), was studied in classification tasks.. In this work, we answer the question: How do we cost-effectively enhance the performance of these models over a continually growing vocabulary space?
First, we investigate continual training of these open-vocabulary models on data with new categories. We consistently find that continual training of these models causes catastrophic forgetting by overwriting past knowledge from updating a limited, poorly understood parameter space. Hence, we investigate how to expand the knowledge of the segmentation model given in-domain data without any training.
Our novel method, kNN-CLIP, simply augments the CLIP segmentation models with a retrieval database that matches images with text descriptions. This database can be updated with new data in a single pass without storing any previous images or requiring training, similar to ACM (Prabhu et al., 2023a). Unlike conventional continual learning methods, kNN-CLIP ensures the model retains knowledge of previously encountered data (zero forgetting), as the distance of a data point to itself in the retrieval database is zero. kNN-CLIP requires a single pass (online), is memory-efficient due to only storing features, and expands its vocabulary with minimal computational resources due to no additional training. The dynamic nature of kNN-CLIP is particularly advantageous for segmentation settings where large volumes of new data are collected daily (Gui et al., 2024), as retraining large foundation models frequently is too expensive.
We demonstrate the effectiveness of kNN-CLIP on state-of-the-art open-vocabulary semantic and panoptic segmentation benchmarks with the FC-CLIP (Yu et al., 2023) model. kNN-CLIP achieves notable performance increases (mIoU) across various challenging datasets: A-847, PC-459, and A-150. Specifically, we see improvements of +2.6, +1.7 and +7.2 points respectively, demonstrating effective segmentation across a continually-growing vocabulary space. This improvement comes with minimal computational and memory overheads and no training cost. As shown in Fig 1, we also qualitatively demonstrate the effectiveness of our method on tasks beyond current benchmarks, such as concept customization and segmenting long-tailed concepts like celebrities, demonstrating the effectiveness of retrieval for tasks considered challenging for CLIP models (Udandarao et al., 2024). Overall, the key contributions of our study include:
  • Dynamically Updating Vocabulary Space: We observe that continual segmentation methods typically underperform compared to zero-shot open vocabulary models. Therefore, we propose adapting the task of continually integrating new classes in continual learning to open vocabulary segmentation models, offering a more feasible solution for continual segmentation.
@@
  • Training-Free Continual Vocabulary Expansion: We introduce a novel technique, kNN-CLIP, that continually expands the vocabulary of image segmentation models beyond the vocabulary space of CLIP without additional training. This method utilizes a continually expanding support set, recasting the problem of concept prediction as an image retrieval task.
@@ 2 Related Work
Retrieval-Augmented Models. In the NLP domain, Retrieval-Augmented Generation (RAG) has been shown as a promising technique for enhancing Large Language Models (LLM) by leveraging external structured data (Guu et al., 2020; Wu et al., 2022; Borgeaud et al., 2022; Liu et al., 2020; Peters et al., 2019; Yu et al., 2021; Izacard et al., 2022; Ram et al., 2023; Yuan et al., 2024). The dynamic nature of RAG facilitates continuous knowledge updates, enabling models to integrate domain-specific information seamlessly. This is particularly beneficial for knowledge-intensive tasks (Lewis et al., 2020; Wang et al., 2022; Khandelwal et al., 2019; Shi et al., 2023; Petroni et al., 2023). Inspired by its success in NLP, researchers are now exploring the application of RAG in computer vision tasks (Chen et al., 2022a; Marino et al., 2021; Roth et al., 2022; Shen et al., 2022; Sheynin et al., 2022; Yasunaga et al., 2022; Blattmann et al., 2022; Chen et al., 2022b; Liu et al., 2023b; Udandarao et al., 2023; Iscen et al., 2023; Prabhu et al., 2023c; Balazevic et al., 2023). For instance, REACT (Liu et al., 2023b) presents a methodology aimed at retrieving relevant knowledge and learning customized visual modules accordingly for specific domains. Meanwhile, SuS-X (Udandarao et al., 2023) and C2C (Prabhu et al., 2023c) introduce frameworks that utilize retrieved support sets for far more accurate CLIP classification to open-set domains. Although these two works show promise in applying retrieval-based methods to visual perception, they did not discuss how to apply such techniques successfully in visual segmentation. Hummingbird (Balazevic et al., 2023) suggests employing a straightforward non-parametric nearest neighbour retrieval as the decoder for dense scene understanding tasks. Moreover, beyond integrating new visual features, RECO (Iscen et al., 2023) demonstrates the efficacy of incorporating text representations by cross-modal fusion alongside original and retrieved embeddings. However, both Hummingbird (Balazevic et al., 2023) and RECO (Iscen et al., 2023) require training to better integrate retrieval augmentation into their pipelines; in contrast, our method demands no training efforts while providing faster adaptability and strong performance.
@@
Open Vocabulary Learning in Image Understanding. Benefiting from advancements in vision-language models (Radford et al., 2021a; Jia et al., 2021), visual models have shown the potential to overcome the constraints of pre-defined closed-set concepts, leading to more flexible open vocabulary image understanding (Li et al., 2023b). This capability is particularly significant for dense prediction tasks (Yu et al., 2023; Rao et al., 2022; Ding et al., 2022; Chen et al., 2023; Karazija et al., 2023; Liang et al., 2023; Li et al., 2022; Xu et al., 2023a; b), where approaches like SimBaseline (Xu et al., 2022b) and OVSeg (Liang et al., 2023) leverage cross-modality supervision from CLIP to align class-agnostic mask proposals with language concepts. Building on these ideas, ODISE (Xu et al., 2023a) and FC-CLIP (Yu et al., 2023) further enhance performance in both semantic and panoptic segmentation by improving visual encoders and mask proposal strategies. Despite these advancements, a critical limitation is the issue of catastrophic forgetting during fine-tuning on concept-restricted dense prediction datasets, which leads to a significant reduction in model vocabulary size (Xu et al., 2022a; Sun et al., 2023a). To mitigate this challenge, training-free methods such as ReCO (Shin et al., 2022) and CaR (Sun et al., 2023a) have been proposed to enable efficient adaptation of CLIP models through recurrent pruning of the label space. While these methods avoid catastrophic forgetting, they still struggle to accommodate continually expanding vocabularies (Wang et al., 2023b; Lüddecke & Ecker, 2022; Xie et al., 2023; Sun et al., 2023b; Seifi et al., 2024). Therefore, in this work, we explore the use of retrieval augmentation to enhance continual open vocabulary image understanding.
@@
Continual Learning for Semantic Segmentation. Unlike open vocabulary semantic segmentation, continual learning does not aim to initially encompass an extremely large vocabulary space but instead focuses on preserving the ability to expand this space continuously. However, continual learning is challenged by issues such as catastrophic forgetting and semantic drift. To address these problems, iCaRL (Rebuffi et al., 2017) suggests replaying the most representative samples during the continual learning stage, and subsequent works have sought to optimize the associated memory burden (Cha et al., 2021; Zhu et al., 2023; Wang et al., 2023a). The use of web images for replay in a similar manner has also been explored (Liu et al., 2023a). Additionally, ALIFE (Oh et al., 2022) introduces a feature-replay scheme to reduce memory requirements. Other research has delved into leveraging memorized features to preserve known knowledge (Yoon et al., 2022; Yu et al., 2020; Wang et al., 2021; Michieli & Zanuttigh, 2021; Douillard et al., 2021). While the use of memorized features shows promise, a significant limitation is that the representation capacity of these features might be insufficient, which can hinder the performance of continual learning (Yuan & Zhao, 2023). Approaches such as employing hyper-class knowledge (Shi et al., 2022) or dynamically updating stored features (Liu et al., 2022; Lin et al., 2022) attempt to mitigate this issue but do not fully circumvent it. Our method, which aims to expand the vocabulary space dynamically, resembles feature-replay methods as we also construct a support set of features. However, our approach differs by operating without the need for additional training, utilizing a support set that stores powerful features learned through self-supervised learning techniques (Caron et al., 2021), with minimal memory requirements.
3 kNN-CLIP: Continually Expanding Retrieval-Augmented Dense Prediction
@@
We introduce a novel, training-free framework for continual dense prediction with expanding vocabularies, applicable to diverse domains and across dense prediction tasks. Inspired by the Retrieval-Augmented Generation (RAG) method used in Large Language Models, our framework leverages a customizable embedding database to incorporate domain-specific knowledge directly at inference, with no additional training.
@@
This approach is applicable to a range of dense prediction tasks including semantic and panoptic image segmentation. We aim to augment existing open-vocabulary segmentation models. These models, including FC-CLIP (Yu et al., 2023) which we use as our baseline, usually have a mask proposer trained to generate class-agnostic masks from extracted features. Then, the image CLIP features are pooled using these proposed masks and used to compute similarities with the text CLIP features of class names for classification. Finally, mask predictions and their corresponding class predictions are combined to give the result segmentation masks. We first evaluate the confidence level of CLIP classification results for each query mask. For queries falling below a given confidence threshold, we extract DINO features and use mask pooling to generate query embeddings. These embeddings are then matched against a vectorized database using a kNN search algorithm based on cosine similarities. For the most similar embeddings found, we create a set of confidence-aware pseudo-logits by stacking similarities across class vectors, guided by the labels of these embeddings. The results from this retrieval process are then combined with the initial CLIP results through a given weighting parameter $\lambda$. We now provide details of our training-free method. We first describe the embedding construction process and detail the query mechanisms.
@@ 3.1 Designing Continually Expanding Embedding Databases
To obtain distinctive vectorized representations for the database images, we employ mask-pooling on features extracted by pretrained encoders from the images as shown in Fig 2. Our input consists of an incoming image ($x_{i}\in\mathcal{R}^{3\times h\times w}$) with its corresponding $k$ class-agnostic masks, which could be semantic, instance dense masks, or bounding box masks ($M_{i}\in\mathcal{R}^{k\times h\times w}$) and class annotations ($C_{i}$) containing the class $\{c{{}_{ij}}\}_{j=1}^{k}$ corresponding to mask $\{m_{ij}\}_{j=1}^{k}=M_{i}$. We extract features of this image using a Vision Transformer (ViT, Dosovitskiy et al. (2021))-based encoder ($\mathcal{E}$), which is pretrained DINOv2 (Oquab et al., 2023) in this case, yielding $h_{i}=\mathcal{E}(x_{i})$. Here, $h_{i}\in\mathcal{R}^{d\times h^{\prime}\times w^{\prime}}$ with $d$ representing the embedding dimension of the feature, and $h^{\prime}$, $w^{\prime}$ represent the dimensions of the feature map. The embedding dimensions $h^{\prime}$, $w^{\prime}$ are determined by image size divided by the patch size of the vision transformer, thus divided by 14 (the patch size of ViT-G) each, and we reduce the space requirements by storing small feature maps rather than input images. The ground-truth masks $M_{i}$ are resized to $M^{\prime}_{i}\in\mathcal{R}^{k\times h^{\prime}\times w^{\prime}}$ to align with the feature map dimensions. By applying mask average pooling along $h^{\prime}$ and $w^{\prime}$, the final embedding sets $\{e_{ij}\}_{j=1}^{k}$ on image $x_{i}$ is obtained by:
$E_{i}=\{e_{ij}=\frac{\sum_{h^{\prime}}\sum_{w^{\prime}}(h_{i}\cdot m^{\prime}_{ij})}{\sum_{h^{\prime}}\sum_{w^{\prime}}m^{\prime}_{ij}}\}_{j=1}^{k},where\ E_{i}\in\mathcal{R}^{k\times d}$ (1)
Then, the vectorized embedding database is constructed by concatenating the embedding sets and class sets from all images, so that $D_{vec}=\{{E_{i},C_{i}}\}_{i=1}^{n}=\{{e_{i^{\prime}},c_{i^{\prime}}}\}_{i^{\prime}=1}^{n^{\prime}}$, if there are $n^{\prime}$ embeddings in total. New embeddings can also be continually added to the database repeating the above process. Note that we do not store any images, enabling low storage costs and allowing deletion of previously seen data.
One appealing characteristic of our framework is the dynamic extendability of the database and the customization of the database construction based on personalized interests. Depending on the purpose specified by the user-defined applications (fine-grained settings such as smart retail, species classification, etc.), the database could be sourced from various domains through manual curation, online web-based collection, or synthetic data sources.